# Flow to the Mode: Mode-Seeking Diffusion Autoencoders for State-of-the-Art Image Tokenization

### v1

@@ 3.4 Sampling
given an initial $x_{1}\sim\mathcal{N}(0,I)$. As an aside, FlowMo can also be used to compute sample log-likelihoods by solving the flow ODE in reverse following flow matching [35].
@@
setting $\rho=1$ reduces to the usual linearly-spaced rectified flow ODE sampler. In the extreme, letting $\rho\rightarrow\infty$ means taking a single large step at $t=1$, which corresponds to regressing $x$ given $c$ [7]. We use $\rho=4$, which corresponds to taking large steps closer to $t=1$ and concentrating the sample towards the mean of $p(x|c)$, while still spending significant sampling FLOPs at the low noise levels, a choice which prior work has shown to be critical for rFID [26]. Our sampler significantly improves both rFID and PSNR.
@@ 3.5 Stage 2: Latent generative modeling
As in other work training autoencoders with discrete latent spaces [16, 52, 60, 37], we verify that our tokenizer can be used to train a high-quality second-stage generative model. We use MaskGiT [8] and our settings for this stage are largely taken from MaskGiT and TiTok [8, 61]. Additional details are in the supplementary material.
@@ 4.1 Main results
Tokenization. For the main task of tokenization, all tokenizers take an image input, encode it to a quantized latent, and then reconstruct the image. All tokenizers are trained on ImageNet-1K [13]. Reconstruction quality is measured in rFID [22], PSNR, SSIM, and LPIPS [62]. We evaluate on the ImageNet-1K validation set, at 256$\times$256 resolution.
The amount of information contained in a tokenizer’s quantized latent code is measured in bits per pixel (BPP). Where $S$ is the latent sequence length and $V$ is the token vocabulary size, the BPP of a tokenizer is computed as $\frac{S\log_{2}(V)}{256^{2}}$. Tokenizers trained at different BPP cannot be compared apples-to-apples since access to more bits will improve performance. Therefore, we train two models in order to match existing state-of-the-art tokenizers. FlowMo-Lo is trained at 0.0703 BPP to match the BPP of OpenMagViT-V2 [37]. FlowMo-Hi is trained at 0.219 BPP to match the BPP of LlamaGen-32 [52]. Both FlowMo-Lo and FlowMo-Hi are exactly the same FlowMo architecture except for the BPP. To match the BPP of the tokenizer against which we compare, we modify the FlowMo token vocabulary size or number of tokens.
We achieve state-of-the-art results at both BPPs in terms of rFID [22], PSNR, and SSIM. The only exception to FlowMo’s strong performance is with respect to the LPIPS metric, for which FlowMo still tends to underperform. Extended visual comparisons are available in the supplementary material and on our website.
@@
Generation. For the task of generation, we train the generative model MaskGiT [8] over the encoded token sequences produced by different tokenizers. We then decode the generated token sequences via the respective tokenizers, and measure the image quality in terms of image generation metrics, namely FID [22], sFID [40], Inception Score [48], and Precision and Recall [33]. The generation metrics are evaluated on the ImageNet-1K benchmark. Despite achieving stronger tokenization, using FlowMo as a tokenizer for a downstream generative model improves some but not all generation metrics. There is an interesting and complicated interplay between tokenizer quality and generation quality, and we hope to improve the results in future work.
Training hyperparameters such as model size, batch size, and training length were identical in Table 2 save the tokenizer. Specifically, we train for 300 epochs atop both tokenizers with batch size 128 and learning rate $1\times 10^{-4}$. Further hyperparameter settings are given in the supplementary material. While we lack the resources to train state-of-the-art generative models, which require e.g. $1,000-1,080$ epochs of training with batch size $1,024-2,048$ [56, 61], our goal here is not to set a state of the art for generation, but to fairly compare tokenizers from the perspective of the latent space they provide for generation.
@@
Table 3: DiTo comparison. Comparison with concurrent work DiTo [10]. Results from the original paper.
Other comparisons. We also conduct a tokenization comparison between FlowMo and DiTo [10], a similar concurrent work, in Table 3. Exclusively for this comparison, we train a tokenizer with a continuous latent space, with LayerNorm as the final encoder layer and the “noise sync” augmentation as proposed in DiTo. We also equalize the overall latent size, using 256 tokens with a token dimension of 16.
Experimental details. All our models are parameterized in $\mu P$ [57] so that hyperparameters can be “$\mu$-transferred” between exploratory and scaled-up configurations. The majority of our experiments were conducted on mixed hardware (A6000, L40S, A100, H100, or H200 GPUs as available) with ViT patch size 8 and hidden dimension 768; this is also the configuration used for ablation studies in Table 4.
Our final experiments (FlowMo-Lo, FlowMo-Hi) were conducted on an 8$\times$×H100 node using a scaled-up model with patch size 4 and hidden dimension increased to 1152 for the decoder only. All other hyperparameters were $\mu-$transferred directly. We pre-train FlowMo-Lo for 100 epochs and FlowMo-Hi for 80 epochs, and then post-train both for approximately 1 epoch. In pre-training, performance in rFID does not saturate. Longer training would require more resources but likely further benefit performance.
We train both FlowMo-Lo and FlowMo-Hi with batch size 128 and learning rate $10^{-4}$. We forcibly normalize the weight matrices in the MLP blocks of the encoder and decoder per step, following EDM2 [30], in order to counteract exploding activations and weight matrices. All models are trained in PyTorch [42] with bfloat16 precision. We use the Adam optimizer [31] with $(\beta_{1},\beta_{2})=(0.9,0.95)$ since we noticed higher $\beta_{2}$ was unstable, possibly due to bfloat16 precision or the long transformer sequence length. We set the encoder learning rate to 00 after $200,000$ training steps. We use an exponential moving average with rate $0.9999$.
@@ 4.2 Analysis
Doubled patch size. In vision transformers [15] and derived architectures [17], “patch size” determines the length of the sequence to which an image is converted, with smaller patch sizes resulting in more computationally intensive models. Prior work on transformer-based diffusion models in pixel space has noted the importance of small patch size [27]. We corroborate this: increasing patch size compromises all metrics.
Encoder trained with MSE. Prior diffusion autoencoder works have explored first training an autoencoder with an MSE- or LPIPS-based regression objective, potentially with an adversarial loss as well, and then using the resulting frozen features as conditioning for a diffusion decoder [2, 26, 41]. Though this style of training may boost PSNR, for best rFID it is essential to train the entire system end-to-end at all noise levels to ensure the latent code contains helpful features for velocity estimation at all noise levels.
No perceptual loss. Prior work in diffusion autoencoders [58] has noted the importance of a perceptual loss on the 1-step denoised prediction of the diffusion decoder. Without a perceptual loss, performance degrades.
FSQ quantization. Compared with our default quantization strategy LFQ [60], we observe that FSQ [39] achieves slightly better final training loss and pairwise metric performance, but at the cost of rFID, which is most critical. We experimented with alternative quantization techniques as well [63, 20], but achieved best performance with LFQ.
Logit-normal noise. The logit-normal noise schedule for rectified flow training [17] assigns 0 probability mass to $t\in\{0,1\}$. Although this is appropriate for $t=0$, where the rectified flow objective becomes singular [38], since our conditioning signal is extremely strong, the estimate of $v(x_{t},t)$ at $t=1$ (pure noise) is critical. Using the usual logit-normal schedule causes PSNR and rFID to collapse. Instead, we use a thick-tailed logit-normal noise schedule: we sample the noise level from a uniform distribution on $[0,1]$ $10\%$ of the time. Without this modification, the velocity estimate at $t=1$ is inaccurate, leading to low PSNR and discoloration. See Figure 7 for an illustrative example.
@@
Figure 7: Noise schedule ablation. Training FlowMo with a logit-normal noise schedule [17] results in discoloration.
@@
No guidance. One advantage of FlowMo over a GAN tokenizer is the ability to leverage techniques such as classifier-free guidance [23] and guidance intervals [34] to improve perceptual quality. Without applying classifier guidance in a limited interval, rFID is worsened.
@@
The primary limitation of FlowMo is inference time. FlowMo requires multiple model forward passes (we use $n=25$ steps in our work) to generate a sample from the decoder given a quantized latent code. By contrast, nearly all other tokenizers are GAN-based and so reconstruct the given image in a single forward pass. Still, there has been considerable work on distilling diffusion models into single- or few-step systems [36, 51], and these techniques are also applicable to FlowMo.
@@ 5 Conclusion
We have presented FlowMo, a transformer-based diffusion autoencoder that achieves state-of-the-art image tokenization by exploiting the key insight that for tokenization, diffusion autoencoder training and sampling should be geared towards seeking perceptually high-quality modes. We have shown state-of-the-art performance on the competitive ImageNet-1K benchmark at 256$\times$256 resolution without using 2D spatially aligned latent codes, adversarial losses, proxy objectives from auxiliary tokenizers, or convolutions, unlike much prior work.
@@ A.1 Stage 1A - Tokenizer pre-training
All hyperparameters are given in Table 6. We found that for FlowMo, using a relatively low commitment loss weight relative to GAN-based tokenizers was beneficial. In fact, bounding the commitment loss by applying a $\mathrm{tanh}$ activation prior to quantization degraded performance. We use token factorization with a $(9,9)$ split following [60, 37], which enables a large vocabulary size without excessive memory usage in the calculation of the entropy loss. We experimented with larger and smaller latent sequence lengths and did not find a significant performance benefit.
Due to limited computational resources, we parameterized all FlowMo models in $\mu P$ and determined the optimal losses, noise schedules, and other hyperparameters in the exploratory “FlowMo (fewer params)” configuration, which was trained on A100, H100, H200, L40S, or A6000 GPUs as available. We then transfer these hyperparameters to the larger and more FLOP-intensive FlowMo-Lo and FlowMo-Hi configurations for final experiments. We train these final configurations for relatively few epochs (100 and 80 respectively) on a single 8×8$\times$8 ×H100 node, and it is likely that further training would continue to improve results.
@@
Num. epochs 40 100 80
$\lambda_{ent}$ 0.1 - -
$\lambda_{commit}$ 0.0025 - -
@@
Total number of parameters (×$(\times 10^{6})$ 517 945 945
@@ A.2 Stage 1B - Tokenizer post-training
MaskBiT [56] noted the effectiveness of a ResNet-style perceptual loss for tokenizer training, and that this loss may have been used in prior work whose tokenizer training recipe was unknown [8]. However, we were unable to use this network effectively in our loss $\mathcal{L}_{\textrm{percep}}$, instead obtaining better results in Stage 1A with the LPIPS-VGG network. Interestingly, [58] also applied a perceptual loss based on LPIPS-VGG on the 1-step denoised prediction of the network.
@@
Table 7: $\mathcal{L}_{\textrm{sample}}$ ablation. We show the result of training with $\mathcal{L}_{\textrm{sample}}$ in the first row. We conduct two sweeps, first to find the optimal loss weight order of magnitude (middle rows) and then to fine-tune the loss weight (bottom rows). Regardless, it cannot match the performance of $\mathcal{L}_{\textrm{sample}}$ Backpropagation through a complete sampling chain is important.
During post-training, due to the computational expense, we reduce the batch size to 64 and correspondingly reduce the learning rate to $5\times 10^{-5}$. In addition to gradient checkpointing, we also use gradient accumulation in this stage. We train Stage 1B for approximately 1 epoch for both FlowMo-Lo and FlowMo-Hi, and apply early stopping to counteract eventual reward hacking [3, 44].
@@ A.3 Stage 2 - Generative model training
To verify the quality of an image tokenizer, it is necessary to ensure it can be used to train a high-quality generative model. As in tokenizer training, we lack resources for matching the computational requirements of state-of-the-art generative models which are trained for e.g., 1000-1080 epochs with batch size 1024-2048 [61, 56]. We use a MaskGiT configuration with hidden size 1024 and 28 layers, with 397M total parameters. training for 300 epochs with learning rate $10^{-4}$ and batch size 128. Both [37] and FlowMo produce 256 tokens of size 18 bits, which we convert to 512 tokens of size 9 bits for training MaskGiT. We use guidance 1.5 for both models, and tune the sampling temperature separately for each model to ensure fair comparison, since we find that models trained atop OpenMagViT-V2 prefer lower temperature sampling, while those trained atop FlowMo prefer high temperature. We use 64 sampling steps for both models with softmax temperature annealing and no guidance decay.
@@ Appendix B Evaluation setup and hyperparameters
For the FlowMo tokenizer, we use a guidance interval [34] of $(0.17,1.02)$ in the EDM2 [30] convention (translating to (0.145, 0.505) in terms of rectified flow noise levels), with classifier-free guidance weight 1.5. For the second stage, we also apply guidance, as is common practice, by shifting the token logits according to the logit difference with and without conditioning.
@@
For all evaluation metrics (rFID, PSNR, etc.) reported in the main paper, we use 50,000 samples unless otherwise noted. We compute the rFID and other reconstruction metrics in Stage 1 against the entire 50,000-example validation set. Generative metrics such as gFID are computed against the full ImageNet-1K training dataset statistics, as is standard practice. The generation results are evaluated using the ADM codebase [14].
Appendix C Rate-distortion-perception tradeoff

### v3

@@ 3.4 Sampling
given an initial $x_{1}\sim\mathcal{N}(0,I)$. As an aside, FlowMo can also be used to compute sample log-likelihoods by solving the flow ODE in reverse following flow matching [flow_matching].
@@
setting $\rho=1$ reduces to the usual linearly-spaced rectified flow ODE sampler. In the extreme, letting $\rho\rightarrow\infty$ means taking a single large step at $t=1$, which corresponds to regressing $x$ given $c$ [genvs]. We use $\rho=4$, which corresponds to taking large steps closer to $t=1$ and concentrating the sample towards the mean of $p(x|c)$, while still spending significant sampling FLOPs at the low noise levels, a choice which prior work has shown to be critical for rFID [highfidelity]. Our sampler significantly improves both rFID and PSNR.
@@ 3.5 Stage 2: Latent generative modeling
As in other work training autoencoders with discrete latent spaces [vqgan, llamagen, magvit_v2, openmagvit_v2], we verify that our tokenizer can be used to train a high-quality second-stage generative model. We use MaskGiT [maskgit] and our settings for this stage are largely taken from MaskGiT and TiTok [maskgit, titok]. Additional details are in the supplementary material.
@@ 4.1 Main results
Tokenization. For the main task of tokenization, all tokenizers take an image input, encode it to a quantized latent, and then reconstruct the image. All tokenizers are trained on ImageNet-1K [imagenet]. Reconstruction quality is measured in rFID [frechet], PSNR, SSIM, and LPIPS [lpips]. We evaluate on the ImageNet-1K validation set, at $\times256$ resolution.
The amount of information contained in a tokenizer’s quantized latent code is measured in bits per pixel (BPP). Where $S$ is the latent sequence length and $V$ is the token vocabulary size, the BPP of a tokenizer is computed as $\frac{S\log_{2}(V)}{256^{2}}$. Tokenizers trained at different BPP cannot be compared apples-to-apples since access to more bits will improve performance. Therefore, we train two models in order to match existing state-of-the-art tokenizers. FlowMo-Lo is trained at 0.0703 BPP to match the BPP of OpenMagViT-V2 [openmagvit_v2]. FlowMo-Hi is trained at 0.219 BPP to match the BPP of LlamaGen-32 [llamagen]. Both FlowMo-Lo and FlowMo-Hi are exactly the same FlowMo architecture except for the BPP. To match the BPP of the tokenizer against which we compare, we modify the FlowMo token vocabulary size or number of tokens.
We achieve state-of-the-art results at both BPPs in terms of rFID [frechet], PSNR, and SSIM. The only exception to FlowMo’s strong performance is with respect to the LPIPS metric, for which FlowMo still tends to underperform. Extended visual comparisons are available in the supplementary material and on our website.
@@
Generation. For the task of generation, we train the generative model MaskGiT [maskgit] over the encoded token sequences produced by different tokenizers. We then decode the generated token sequences via the respective tokenizers, and measure the image quality in terms of image generation metrics, namely FID [frechet], sFID [sfid], Inception Score [inception_score], and Precision and Recall [precrec]. The generation metrics are evaluated on the ImageNet-1K benchmark. Despite achieving stronger tokenization, using FlowMo as a tokenizer for a downstream generative model improves some but not all generation metrics. There is an interesting and complicated interplay between tokenizer quality and generation quality, and we hope to improve the results in future work.
Training hyperparameters such as model size, batch size, and training length were identical in Table 2 save the tokenizer. Specifically, we train for 300 epochs atop both tokenizers with batch size 128 and learning rate $1\times 10^{-4}$. Further hyperparameter settings are given in the supplementary material. While we lack the resources to train state-of-the-art generative models, which require e.g. 1,0001,0801,000-1,080 epochs of training with batch size 1,0242,0481,024-2,048 [maskbit, titok], our goal here is not to set a state of the art for generation, but to fairly compare tokenizers from the perspective of the latent space they provide for generation.
@@
Table 3: DiTo comparison. Comparison with concurrent work DiTo [dito]. Results from the original paper.
Other comparisons. We also conduct a tokenization comparison between FlowMo and DiTo [dito], a similar concurrent work, in Table 3. Exclusively for this comparison, we train a tokenizer with a continuous latent space, with LayerNorm as the final encoder layer and the “noise sync” augmentation as proposed in DiTo. We also equalize the overall latent size, using 256 tokens with a token dimension of 16.
Experimental details. All our models are parameterized in $\mu P$ [mup] so that hyperparameters can be $\mu$-transferred” between exploratory and scaled-up configurations. The majority of our experiments were conducted on mixed hardware (A6000, L40S, A100, H100, or H200 GPUs as available) with ViT patch size 8 and hidden dimension 768; this is also the configuration used for ablation studies in Table 4.
[…]
We train both FlowMo-Lo and FlowMo-Hi with batch size 128 and learning rate $10^{-4}$. We forcibly normalize the weight matrices in the MLP blocks of the encoder and decoder per step, following EDM2 [edm2], in order to counteract exploding activations and weight matrices. All models are trained in PyTorch [pytorch] with bfloat16 precision. We use the Adam optimizer [adam] with $(\beta_{1},\beta_{2})=(0.9,0.95)$ since we noticed higher $\beta_{2}$ was unstable, possibly due to bfloat16 precision or the long transformer sequence length. We set the encoder learning rate to 0 after $200,000$ training steps. We use an exponential moving average with rate $0.9999$.
@@ 4.2 Analysis
Doubled patch size. In vision transformers [vit] and derived architectures [sd3], “patch size” determines the length of the sequence to which an image is converted, with smaller patch sizes resulting in more computationally intensive models. Prior work on transformer-based diffusion models in pixel space has noted the importance of small patch size [simpler_diffusion]. We corroborate this: increasing patch size compromises all metrics.
Encoder trained with MSE. Prior diffusion autoencoder works have explored first training an autoencoder with an MSE- or LPIPS-based regression objective, potentially with an adversarial loss as well, and then using the resulting frozen features as conditioning for a diffusion decoder [swycc, highfidelity, dalle3]. Though this style of training may boost PSNR, for best rFID it is essential to train the entire system end-to-end at all noise levels to ensure the latent code contains helpful features for velocity estimation at all noise levels.
No perceptual loss. Prior work in diffusion autoencoders [yangmandt] has noted the importance of a perceptual loss on the 1-step denoised prediction of the diffusion decoder. Without a perceptual loss, performance degrades.
FSQ quantization. Compared with our default quantization strategy LFQ [magvit_v2], we observe that FSQ [fsq] achieves slightly better final training loss and pairwise metric performance, but at the cost of rFID, which is most critical. We experimented with alternative quantization techniques as well [bsq, rqvae], but achieved best performance with LFQ.
Logit-normal noise. The logit-normal noise schedule for rectified flow training [sd3] assigns 0 probability mass to $t\in\{0,1\}$. Since our conditioning signal is extremely strong, the estimate of $v(x_{t},t)$ at $t=1$ (pure noise) is critical. Using the usual logit-normal schedule causes PSNR and rFID to collapse. Instead, we use a thick-tailed logit-normal noise schedule: we sample the noise level from a uniform distribution on $[0,1]$ $10\%$ of the time. Without this modification, the result is low PSNR and discoloration. See Figure 7 for an illustrative example.
@@
Figure 7: Noise schedule ablation. Training FlowMo with a logit-normal noise schedule [sd3] results in discoloration.
@@
No guidance. One advantage of FlowMo over a GAN tokenizer is the ability to leverage techniques such as classifier-free guidance [cfg] and guidance intervals [cfg_interval] to improve perceptual quality. Without applying classifier guidance in a limited interval, rFID is worsened. We enable guidance by dropping out the quantized latent code $c$ $10\%$ of the time during tokenizer training. We do not leverage any class information in the tokenizer.
@@
The primary limitation of FlowMo is inference time. FlowMo requires multiple model forward passes (we use $n=25$ steps in our work) to generate a sample from the decoder given a quantized latent code.
@@ 5 Conclusion
We have presented FlowMo, a transformer-based diffusion autoencoder that achieves state-of-the-art image tokenization. We have shown state-of-the-art performance on the competitive ImageNet-1K benchmark at $\times256$ resolution without using 2D spatially aligned latent codes, adversarial losses, proxy objectives from auxiliary tokenizers, or convolutions, unlike much prior work.
Acknowledgments. This work is in part supported by ONR MURI N00014-22-1-2740.
[…]
@@ A.1 Stage 1A - Tokenizer pre-training
All hyperparameters are given in Table 6. We found that for FlowMo, using a relatively low commitment loss weight relative to GAN-based tokenizers was beneficial. In fact, bounding the commitment loss by applying a $\mathrm{tanh}$ activation prior to quantization degraded performance. We use token factorization following [magvit_v2, openmagvit_v2], which enables a large vocabulary size without excessive memory usage in the calculation of the entropy loss.
Due to limited computational resources, we parameterized all FlowMo models in $\mu P$ and determined the optimal losses, noise schedules, and other hyperparameters in the exploratory “FlowMo (fewer params)” configuration, which was trained on A100, H100, H200, L40S, or A6000 GPUs as available. We then transfer these hyperparameters to the larger and more FLOP-intensive FlowMo-Lo and FlowMo-Hi configurations for final experiments. We train these final configurations for relatively few epochs (130 and 80 respectively) on a single $8\timesH$100 node, and it is likely that further training would continue to improve results.
@@
Num. epochs 40 130 80
$\lambda_{ent}$ 0.0025 - -
$\lambda_{commit}$ 0.000625 - -
@@
Total number of parameters $(\times 10^{6})$ 517 945 945
@@ A.2 Stage 1B - Tokenizer post-training
MaskBiT [maskbit] noted the effectiveness of a ResNet-style perceptual loss for tokenizer training, and that this loss may have been used in prior work whose tokenizer training recipe was unknown [maskgit]. However, we were unable to use this network effectively in our loss $\mathcal{L}_{\textrm{percep}}$, instead obtaining better results in Stage 1A with the LPIPS-VGG network. Interestingly, [yangmandt] also applied a perceptual loss based on LPIPS-VGG on the 1-step denoised prediction of the network.
@@
Table 7: $\mathcal{L}_{\textrm{sample}}$ ablation. We show the result of training with $\mathcal{L}_{\textrm{sample}}$ in the first row. We then compare with $\mathcal{L}_{\textrm{perc}}$ at various weights. We conduct two sweeps, first to find the optimal loss weight order of magnitude (middle rows) and then to fine-tune the loss weight (bottom rows). Regardless, it cannot match the performance of $\mathcal{L}_{\textrm{sample}}$ Backpropagation through a complete sampling chain is important.
During post-training, due to the computational expense, we reduce the batch size to 64 and correspondingly reduce the learning rate to $5\times 10^{-5}$. In addition to gradient checkpointing, we also use gradient accumulation in this stage. We train Stage 1B for approximately 1 epoch for both FlowMo-Lo and FlowMo-Hi, and apply early stopping to counteract eventual reward hacking [ddpo, alignprop].
@@ A.3 Stage 2 - Generative model training
To verify the quality of an image tokenizer, it is necessary to ensure it can be used to train a high-quality generative model. As in tokenizer training, we lack resources for matching the computational requirements of state-of-the-art generative models which are trained for e.g., 1000-1080 epochs with batch size 1024-2048 [titok, maskbit]. We use a MaskGiT configuration with hidden size 1024 and 28 layers, with 397M total parameters. training for 300 epochs with learning rate $10^{-4}$ and batch size 128. Both [openmagvit_v2] and FlowMo produce 256 tokens of size 18 bits, which we convert to 512 tokens of size 9 bits for training MaskGiT. We use guidance 1.5 for both models, and tune the sampling temperature separately for each model to ensure fair comparison, since we find that models trained atop OpenMagViT-V2 prefer lower temperature sampling, while those trained atop FlowMo prefer high temperature. We use 64 sampling steps for both models with softmax temperature annealing and no guidance decay.
@@ Appendix B Evaluation setup and hyperparameters
For the FlowMo tokenizer, we use a guidance interval [cfg_interval] of $(0.17,1.02)$ in the EDM2 [edm2] convention (translating to (0.145, 0.505) in terms of rectified flow noise levels), with classifier-free guidance weight 1.5. For the second stage, we also apply guidance, as is common practice, by shifting the token logits according to the logit difference with and without conditioning.
@@
For all evaluation metrics (rFID, PSNR, etc.) reported in the main paper, we use 50,000 samples unless otherwise noted. We compute the rFID and other reconstruction metrics in Stage 1 against the entire 50,000-example validation set. Generative metrics such as gFID are computed against the full ImageNet-1K training dataset statistics, as is standard practice. The generation results are evaluated using the ADM codebase [adm].
Appendix C Additional Experiments
In this section, we provide some additional analyses and experiments to better understand and illustrate the capabilities of FlowMo.
C.1 Rate-distortion-perception tradeoff